Name the property under multiplication used in each of the following. 
-4/5 × 1 = 1 × (-4/5) = -4/5

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

-4/5×1=1×(-4/5)=-4/5
Here 1 is the multiplicative identity.

​Thanks